# ButterflyQuant: Ultra-low-bit LLM Quantization through Learnable Orthogonal Butterfly Transforms

### v2

@@ 3.3 Composite Butterfly Transforms for Non-Power-of-2 Dimensions
While butterfly transforms naturally handle power-of-2 dimensions through their recursive structure, many practical LLMs use dimensions that are not powers of 2. For instance, LLaMA-2-13B uses dimension 5120 = 40 × 128, where 40 is not a power of 2. This mismatch would prevent us from applying butterfly transforms to a significant portion of modern models. We address this challenge using composite transforms based on Kronecker products [Loan and Pitsianis, 1993, Lathauwer et al., 2000], which allow us to combine smaller orthogonal transforms while maintaining computational efficiency.
@@ 3.3.1 Kronecker Product Formulation
The use of Kronecker products for efficient transformations in quantization has been explored in recent work. FlatQuant [Liu et al., 2024c] employs Kronecker decomposition $\mathbf{P}=\mathbf{P}_{1}\otimes\mathbf{P}_{2}$ to reduce the computational and memory overhead of their affine transformations, where $\mathbf{P}_{1}\in\mathbb{R}^{n_{1}\times n_{1}}$ and $\mathbf{P}_{2}\in\mathbb{R}^{n_{2}\times n_{2}}$ with $n=n_{1}n_{2}$. Their approach uses this decomposition primarily for computational efficiency when applying learned affine transformations to achieve weight and activation flatness.
@@
where $\mathbf{Q}_{1}\in\mathbb{R}^{d_{1}\times d_{1}}$ and $\mathbf{Q}_{2}\in\mathbb{R}^{d_{2}\times d_{2}}$ are orthogonal matrices.
@@
$(\mathbf{Q}_{1}\otimes\mathbf{Q}_{2})^{T}(\mathbf{Q}_{1}\otimes\mathbf{Q}_{2})=(\mathbf{Q}_{1}^{T}\mathbf{Q}_{1})\otimes(\mathbf{Q}_{2}^{T}\mathbf{Q}_{2})=\mathbf{I}_{d_{1}}\otimes\mathbf{I}_{d_{2}}=\mathbf{I}_{d}$ (14)
@@ 3.3.2 Concrete Example for d = 5120
3.4 Loss Function
$\mathcal{L}_{\text{recon}}=\|\mathbf{W}\mathbf{x}-\text{Dequant}(\text{Quant}(\mathbf{W}\mathbf{B}^{T}))\cdot\text{Dequant}(\text{Quant}(\mathbf{B}\mathbf{x}))\|_{2}^{2}$ (17)
and $\text{Dequant}(\cdot)$ is the corresponding dequantization: $\text{Dequant}(q)=s\cdot q$, where $s=\max(|x|)/(2^{b-1}-1)$ is the scale factor computed per tensor or per group. This objective, similar to block-wise reconstruction in GPTQ [Frantar et al., 2023] and learnable quantization methods [Shao et al., 2023], directly minimizes the quantization-induced error while maintaining computational invariance through the butterfly transform.
3.4.1 Uniformity Regularization
While traditional quantization methods apply uniformity regularization to weight distributions [Park et al., 2017, Baskin et al., 2021], we specifically target activations for two key reasons: (1) Theoretical justification from information theory: uniform quantization achieves maximum entropy for a given bit-width, and applying this to activations ensures optimal information preservation through the quantized layer [Cover and Thomas, 2006]; (2) The rotation-quantization duality $\mathbf{y}=\text{Q}(\mathbf{W}\mathbf{B}^{T})\cdot\text{Q}(\mathbf{B}\mathbf{x})$ means that uniformizing activations through $\mathbf{B}$ simultaneously improves both weight and activation quantization.
3.5 Mathematical Properties
Theorem 2 (Expressive Power of Butterfly Transforms).
Butterfly transforms with $\log_{2}n$ layers can exactly represent the Hadamard, Discrete Fourier Transform (DFT), and Discrete Cosine Transform (DCT) matrices, and can approximate any orthogonal matrix to arbitrary precision with sufficient parameters [Cooley and Tukey, 1965, Dao et al., 2022b]. (Proof sketch in Section B)
3.5.1 Gradient Flow Through Butterfly Layers
4 Experiments
Table 1: Comprehensive evaluation of 2-bit weight quantization (W2A16) on LLaMA-2 models. Some results are from the LLM quantization benchmarking paper [Zhao et al., 2025].
@@ 4.1 Experimental Setup
We evaluate ButterflyQuant on LLaMA-2-7B and LLaMA-2-13B [Touvron et al., 2023], using WikiText-2 [Merity et al., 2017] and C4 [Raffel et al., 2020] for perplexity evaluation, plus six zero-shot reasoning tasks: WinoGrande [Sakaguchi et al., 2020], PIQA [Bisk et al., 2020], HellaSwag [Zellers et al., 2019], ARC [Clark et al., 2018], and MMLU [Hendrycks et al., 2021]. We compare against GPTQ [Frantar et al., 2023], AWQ [Lin et al., 2024a], OmniQuant [Shao et al., 2023], QuIP [Chee et al., 2023] and SpinQuant [Liu et al., 2024a] under aggressive 2-bit weight quantization (2-bit weights with 16-bit activations, W2A16). More experiments on other LLM architectures can be found in Appendix.
@@ 4.2 Main Results
Table 1 demonstrates ButterflyQuant’s superiority in low-bit quantization across all metrics. We achieve 2.4× lower perplexity than the best baseline (15.40 vs 36.77 for GPTQ on WikiText-2 7B, 10.24 vs 13.75 for QuIP on 13B) while AWQ fails catastrophically with perplexity exceeding $10^{5}$. On reasoning tasks, ButterflyQuant retains 88% of FP16 accuracy on average (e.g., 62.27% vs 69.06% FP16 on WinoGrande, 68.97% vs 78.07% FP16 on PIQA) while baselines retain only 65-73% of FP16 performance, with consistent improvements across both model scales validating the effectiveness of our approach.
@@ 4.3 Ablation Studies
(a) Impact of initialization strategy on final perplexity.
(b) Convergence analysis showing 86% improvement within 200 steps.
Figure 3: Ablation Studies.
Ablations validate our design choices. Identity initialization (15.428 perplexity) outperforms Hadamard (16.452) and random orthogonal (17.823) initialization, enabling gradual rotation learning through small incremental adjustments (Figure 3(a)). Training converges within 500 steps with 86% of the improvement achieved in just 200 steps (Figure 3(b)), confirming lightweight optimization.
@@ 5 Conclusion
ButterflyQuant demonstrates that continuous parameterization of orthogonal transforms fundamentally changes what is achievable in extreme quantization. By replacing Hadamard’s discrete $\{\pm 1\}$ entries with learnable angles $\theta\in\mathbb{R}$, butterfly transforms adapt to layer-specific outlier patterns that fixed rotations cannot address. This simple but powerful insight enables practical 2-bit quantization with 2.4× lower perplexity than state-of-the-art methods while retaining 88% of FP16 accuracy on average. Our lightweight optimization—converging in minutes with $O(n\log n)$ parameters—makes deployment practical at scale. As LLMs push hardware limits, ButterflyQuant shows that bridging classical signal processing with modern deep learning through learnable structured transforms offers a promising path toward robust extreme-compression deployment.
Appendix A Supplemental Materials
A.1 Implementation Details
Hardware. All experiments use a single NVIDIA H100 GPU with PyTorch 2.2.1 [Paszke et al., 2019]. Following established protocols [Ashkboos et al., 2024, Frantar et al., 2023], we use 128 calibration samples and 2048-token sequences for evaluation.
Appendix B Butterfly Transform Details
B.1 Example: 4×4 Hadamard
For $n=4$, the Hadamard matrix is:
$\mathbf{H}_{4}=\frac{1}{2}\begin{bmatrix}1&1&1&1\\ 1&-1&1&-1\\ 1&1&-1&-1\\ 1&-1&-1&1\end{bmatrix}$
This can be factorized as:
$\mathbf{H}_{4}=\frac{1}{2}\cdot\mathbf{B}_{1}\cdot\mathbf{B}_{2}\cdot\mathbf{P}$
where:
**Layer 1** (pairs (0,1) and (2,3)):
$\mathbf{B}_{1}=\begin{bmatrix}\cos\frac{\pi}{4}&-\sin\frac{\pi}{4}&0&0\\ \sin\frac{\pi}{4}&\cos\frac{\pi}{4}&0&0\\ 0&0&\cos\frac{\pi}{4}&-\sin\frac{\pi}{4}\\ 0&0&\sin\frac{\pi}{4}&\cos\frac{\pi}{4}\end{bmatrix}$
**Layer 2** (pairs (0,2) and (1,3) after permutation):
$\mathbf{B}_{2}=\mathbf{P}_{2}^{T}\begin{bmatrix}\cos\frac{\pi}{4}&0&-\sin\frac{\pi}{4}&0\\ 0&\cos\frac{\pi}{4}&0&-\sin\frac{\pi}{4}\\ \sin\frac{\pi}{4}&0&\cos\frac{\pi}{4}&0\\ 0&\sin\frac{\pi}{4}&0&\cos\frac{\pi}{4}\end{bmatrix}\mathbf{P}_{2}$
B.2 Concrete Example: Butterfly Matrix Structure for $n=8$
To illustrate the butterfly structure concretely, consider an 8-dimensional transform with $\log_{2}8=3$ layers. Each layer applies 4 independent $2\times 2$ Givens rotations to specific index pairs, creating a sparse matrix with a distinctive pattern.
Layer 1 (Stride 1): Pairs adjacent indices $(0,1),(2,3),(4,5),(6,7)$:
$\mathbf{B}_{1}=\begin{bmatrix}\cos\theta_{1,1}&-\sin\theta_{1,1}&0&0&0&0&0&0\\ \sin\theta_{1,1}&\cos\theta_{1,1}&0&0&0&0&0&0\\ 0&0&\cos\theta_{1,2}&-\sin\theta_{1,2}&0&0&0&0\\ 0&0&\sin\theta_{1,2}&\cos\theta_{1,2}&0&0&0&0\\ 0&0&0&0&\cos\theta_{1,3}&-\sin\theta_{1,3}&0&0\\ 0&0&0&0&\sin\theta_{1,3}&\cos\theta_{1,3}&0&0\\ 0&0&0&0&0&0&\cos\theta_{1,4}&-\sin\theta_{1,4}\\ 0&0&0&0&0&0&\sin\theta_{1,4}&\cos\theta_{1,4}\end{bmatrix}$ (21)
Layer 2 (Stride 2): After permutation, pairs indices with stride 2: $(0,2),(1,3),(4,6),(5,7)$:
$\mathbf{B}_{2}=\begin{bmatrix}\cos\theta_{2,1}&0&-\sin\theta_{2,1}&0&0&0&0&0\\ 0&\cos\theta_{2,2}&0&-\sin\theta_{2,2}&0&0&0&0\\ \sin\theta_{2,1}&0&\cos\theta_{2,1}&0&0&0&0&0\\ 0&\sin\theta_{2,2}&0&\cos\theta_{2,2}&0&0&0&0\\ 0&0&0&0&\cos\theta_{2,3}&0&-\sin\theta_{2,3}&0\\ 0&0&0&0&0&\cos\theta_{2,4}&0&-\sin\theta_{2,4}\\ 0&0&0&0&\sin\theta_{2,3}&0&\cos\theta_{2,3}&0\\ 0&0&0&0&0&\sin\theta_{2,4}&0&\cos\theta_{2,4}\end{bmatrix}$ (22)
Note the “crossing” pattern: elements at positions $(0,2)$ and $(2,0)$ are now coupled, creating the characteristic butterfly connections.
Layer 3 (Stride 4): Pairs indices with stride 4: $(0,4),(1,5),(2,6),(3,7)$:
$\mathbf{B}_{3}=\begin{bmatrix}\cos\theta_{3,1}&0&0&0&-\sin\theta_{3,1}&0&0&0\\ 0&\cos\theta_{3,2}&0&0&0&-\sin\theta_{3,2}&0&0\\ 0&0&\cos\theta_{3,3}&0&0&0&-\sin\theta_{3,3}&0\\ 0&0&0&\cos\theta_{3,4}&0&0&0&-\sin\theta_{3,4}\\ \sin\theta_{3,1}&0&0&0&\cos\theta_{3,1}&0&0&0\\ 0&\sin\theta_{3,2}&0&0&0&\cos\theta_{3,2}&0&0\\ 0&0&\sin\theta_{3,3}&0&0&0&\cos\theta_{3,3}&0\\ 0&0&0&\sin\theta_{3,4}&0&0&0&\cos\theta_{3,4}\end{bmatrix}$ (23)
B.3 The Butterfly Pattern Visualization
The name “butterfly” comes from the crossing pattern of connections when visualized as a computational graph. For an 8-point transform with 3 layers:
Input:  0    1    2    3    4    5    6    7
Layer 1 (adjacent pairs):
        0----1    2----3    4----5    6----7
Layer 2 (stride 2):
        0---------2         4---------6
             X                   X
        1---------3         5---------7
Layer 3 (stride 4):
        0-------------------4
        1-------------------5
        2-------------------6
        3-------------------7
Output: 0    1    2    3    4    5    6    7
The crossing patterns (marked with X) in Layer 2 create the characteristic “butterfly wings” shape. Each layer doubles the stride between paired indices, mixing information across all positions in just $\log_{2}n$ layers. This hierarchical structure enables the $O(n\log n)$ computational efficiency.
B.4 Permutation Structure and Block Decomposition
The permutation matrices $\mathbf{P}_{i}$ implement the bit-reversal permutation pattern from the FFT algorithm. After permutation, the rotation matrix becomes block-diagonal:
$\mathbf{P}_{2}^{T}\mathbf{B}_{2}\mathbf{P}_{2}=\begin{bmatrix}\mathbf{G}(\theta_{2,1})&\mathbf{0}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{G}(\theta_{2,2})&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{G}(\theta_{2,3})&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{G}(\theta_{2,4})\end{bmatrix}$ (24)
B.5 Alternative Factorizations for Non-Power-of-2 Dimensions
The choice of factorization affects both expressiveness and efficiency. For $d=5120$, possible factorizations include:
  • $5120=80\times 64$: Both factors closer to powers of 2
  • $5120=20\times 256$: Larger power-of-2 component
  • $5120=5\times 1024$: Minimal non-power-of-2 component
  • $5120=40\times 128$: Balanced factorization (our choice)
Each factorization offers different trade-offs:
  • $80\times 64$: More uniform but requires a composite butterfly for both factors
  • $20\times 256$: Efficient $256=2^{8}$ component but small first factor limits expressiveness
  • $5\times 1024$: Maximizes power-of-2 efficiency but $5\times 5$ is too restrictive
  • $40\times 128$: Balances parameter count (1,228) with expressiveness
Empirical evaluation shows the $40\times 128$ factorization achieves the best quantization performance while maintaining computational efficiency.
B.6 Mathematical Properties
Theorem 3 (Expressive Power of Butterfly Transforms).
Butterfly transforms with $O(\log n)$ layers can efficiently approximate structured orthogonal matrices and exactly represent important transforms, including Hadamard, DFT, and DCT matrices.
Proof Sketch.
While butterfly transforms with $\frac{n\log n}{2}$ parameters cannot represent arbitrary $n\times n$ orthogonal matrices (which have $\frac{n(n-1)}{2}$ degrees of freedom), they form a universal building block for structured matrices—those admitting fast $O(n\log n)$ algorithms [Dao et al., 2019]. Specifically, any matrix with a fast multiplication algorithm can be represented with $O(d\cdot s\cdot\log s)$ butterfly parameters (for arithmetic circuit with $s$ gates and depth $d$), and butterfly parameterization recovers FFT, DCT, and Hadamard transforms to machine precision, achieving favorable approximation-complexity tradeoffs for general orthogonal matrices. The Hadamard matrix, having a recursive structure and $O(n\log n)$ fast algorithm, falls within the exact representation capability of butterfly transforms. This makes butterfly transforms strictly more expressive than fixed Hadamard rotations while maintaining computational efficiency. ∎
B.7 Method Characteristics and Efficiency
Table 2: Method characteristics: training requirements and additional parameters.
Method Calibration Training Time Extra Params Orthogonal
GPTQ 128 samples Minutes None $\times$
AWQ 128 samples Minutes Scales only $\times$
SmoothQuant 512 samples Minutes Scales only $\times$
OmniQuant 128 samples Hours Affine params $\times$
QuaRot None None None $\checkmark$ (Fixed)
SpinQuant 1024 samples Hours $O(n^{2})$ $\checkmark$ (Learned)
ButterflyQuant 128 samples 5-10 min $O(n\log n)$ $\checkmark$ (Learned)
Table 2 highlights ButterflyQuant’s practical advantages: it combines the theoretical guarantees of orthogonal methods with efficient learning, requiring only minutes of optimization compared to hours for SpinQuant, while using exponentially fewer parameters.

### v3

@@ 3.3 Composite Butterfly Transforms for Non-Power-of-2 Dimensions
While butterfly transforms naturally handle power-of-2 dimensions through their recursive structure, many practical LLMs use dimensions that are not powers of 2. For instance, LLaMA-2-13B uses dimension 5120 = 40 × 128, where 40 is not a power of 2. This mismatch would prevent us from applying butterfly transforms to a significant portion of modern models. We address this challenge using composite transforms based on Kronecker products (Loan and Pitsianis, 1993; Lathauwer et al., 2000), which allow us to combine smaller orthogonal transforms while maintaining computational efficiency.
@@ 3.3.1 Kronecker Product Formulation
The use of Kronecker products for efficient transformations in quantization has been explored in recent work. FlatQuant (Sun et al., 2024b) employs Kronecker decomposition $\mathbf{P}=\mathbf{P}_{1}\otimes\mathbf{P}_{2}$ to reduce the computational and memory overhead of their affine transformations, where $\mathbf{P}_{1}\in^{n_{1}\times n_{1}}$ and $\mathbf{P}_{2}\in^{n_{2}\times n_{2}}$ with $n=n_{1}n_{2}$. Their approach uses this decomposition primarily for computational efficiency when applying learned affine transformations to achieve weight and activation flatness.
@@
where $\mathbf{Q}_{1}\in^{d_{1}\times d_{1}}$ and $\mathbf{Q}_{2}\in^{d_{2}\times d_{2}}$ are orthogonal matrices.
@@
$(\mathbf{Q}_{1}\otimes\mathbf{Q}_{2})^{T}(\mathbf{Q}_{1}\otimes\mathbf{Q}_{2})$ $=(\mathbf{Q}_{1}^{T}\mathbf{Q}_{1})\otimes(\mathbf{Q}_{2}^{T}\mathbf{Q}_{2})$ (14)
$=\mathbf{I}_{d_{1}}\otimes\mathbf{I}_{d_{2}}$ $=\mathbf{I}_{d}$
@@ 3.3.2 Concrete Example for d = 5120
Table 1: Comprehensive evaluation of 2-bit weight quantization (W2A16) on LLaMA-2 models. Some results are from the LLM quantization benchmarking paper (Zhao et al., 2025).
@@
3.4 Loss Function
$\mathcal{L}_{\text{recon}}$ $=$ (17)
$\|\mathbf{W}\mathbf{x}-\text{Dequant}(\text{Quant}(\mathbf{W}\mathbf{B}^{T}))$ $\cdot\text{Dequant}(\text{Quant}(\mathbf{B}\mathbf{x}))\|_{2}^{2}$
and $\text{Dequant}(\cdot)$ is the corresponding dequantization: $\text{Dequant}(q)=s\cdot q$, where $s=\max(|x|)/(2^{b-1}-1)$ is the scale factor computed per tensor or per group. This objective, similar to block-wise reconstruction in GPTQ (Frantar et al., 2023) and learnable quantization methods (Shao et al., 2023), directly minimizes the quantization-induced error while maintaining computational invariance through the butterfly transform.
3.4.1 Uniformity Regularization
While traditional quantization methods apply uniformity regularization to weight distributions (Park et al., 2017; Baskin et al., 2021), we specifically target activations for two key reasons: (1) Theoretical justification from information theory: uniform quantization achieves maximum entropy for a given bit-width, and applying this to activations ensures optimal information preservation through the quantized layer (Cover, 1999); (2) The rotation-quantization duality $\mathbf{y}=\text{Q}(\mathbf{W}\mathbf{B}^{T})\cdot\text{Q}(\mathbf{B}\mathbf{x})$ means that uniformizing activations through $\mathbf{B}$ simultaneously improves both weight and activation quantization.
3.5 Mathematical Properties
Theorem 3.2 (Expressive Power of Butterfly Transforms).
Butterfly transforms with $\log_{2}n$ layers can exactly represent important structured orthogonal matrices—including the Hadamard, Discrete Fourier Transform (DFT), and Discrete Cosine Transform (DCT) matrices—and form a universal building block for any structured matrix admitting an $O(n\log n)$ fast algorithm (Cooley and Tukey, 1965; Dao et al., 2022a). While butterfly transforms with $\frac{n\log n}{2}$ parameters cannot represent arbitrary $n\times n$ orthogonal matrices (which require $\frac{n(n-1)}{2}$ degrees of freedom), they achieve favorable approximation-complexity tradeoffs for structured matrices encountered in practice. (Proof sketch in Section 7)
3.5.1 Gradient Flow Through Butterfly Layers
Figure 3: Impact of initialization strategy on final perplexity.
4 Experiments
4.1 Implementation Details
Hardware. All experiments use a single NVIDIA H100 GPU with PyTorch 2.2.1 (Paszke et al., 2019). Following established protocols (Ashkboos et al., 2024; Frantar et al., 2023), we use 128 calibration samples and 2048-token sequences for evaluation.
Figure 4: Convergence analysis showing 86% improvement within 200 steps.
@@ 4.2 Experimental Setup
We evaluate ButterflyQuant on LLaMA-2-7B and LLaMA-2-13B (Touvron et al., 2023), using WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) for perplexity evaluation, plus six zero-shot reasoning tasks: WinoGrande (Sakaguchi et al., 2020), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), ARC (Clark et al., 2018), and MMLU (Hendrycks et al., 2021). We compare against GPTQ (Frantar et al., 2023), AWQ (Lin et al., 2024b), OmniQuant (Shao et al., 2023), QuIP (Chee et al., 2023) and SpinQuant (Liu et al., 2024c) under aggressive 2-bit weight quantization (2-bit weights with 16-bit activations, W2A16). More experiments on other LLM architectures can be found in Appendix.
@@ 4.3 Main Results
Table 1 demonstrates ButterflyQuant’s consistent superiority in low-bit quantization across all metrics. Compared to the strongest baselines, ButterflyQuant reduces WikiText-2 perplexity from 16.43 (SpinQuant) to 15.40 on LLaMA-2-7B (6.3% relative improvement) and from 13.75 (QuIP) to 10.24 on LLaMA-2-13B (25.5% improvement), approaching FP16 quality (4.88). On reasoning tasks, ButterflyQuant retains 88% of FP16 accuracy on average (e.g., 62.27% vs 69.06% FP16 on WinoGrande, 68.97% vs 78.07% FP16 on PIQA) while baselines retain only 65–73% of FP16 performance, with consistent improvements across both model scales validating the effectiveness of layer-adaptive learnable rotations.
@@ 4.4 Ablation Studies
Ablations validate our design choices. Identity initialization (15.428 perplexity) outperforms Hadamard (16.452) and random orthogonal (17.823) initialization, enabling gradual rotation learning through small incremental adjustments (Figure 3). Training converges within 500 steps with 86% of the improvement achieved in just 200 steps (Figure 4), confirming lightweight optimization.
@@
Inference Efficiency.
A practical concern is whether replacing fixed Hadamard rotations with learned butterfly transforms introduces inference overhead. Both transforms share an identical computational graph: $\log_{2}n$ sequential stages, each applying $n/2$ independent $2\times 2$ operations with the same butterfly memory access pattern (strided reads and writes with stride $2^{i-1}$ at stage $i$). The asymptotic complexity is $O(n\log n)$ in both cases. The key difference is arithmetic: FHT stages perform additions and subtractions only (since all entries are $\pm 1$), whereas our butterfly stages require two fused multiply-add (FMA) operations per pair using pre-computed $\cos\theta_{i,j}$ and $\sin\theta_{i,j}$ twiddle factors stored in a lookup table of size $\frac{n\log_{2}n}{2}$. On modern GPUs, butterfly-structured transforms are typically memory-bandwidth bound rather than compute-bound: the dominant cost is the $\log_{2}n$ passes over the $n$-element vector, not the ALU operations within each pass (Dao et al., 2022b). Since ButterflyQuant preserves the same number of passes, the same access pattern, and the same parallelism as FHT, the wall-clock overhead of the additional multiplications is expected to be modest. Concretely, existing CUDA kernels for fast Hadamard transforms (e.g., from QuaRot (Ashkboos et al., 2024)) can be adapted to ButterflyQuant by replacing the fixed $\pm 1$ coefficients with pre-computed twiddle factors, requiring only minor kernel modifications rather than a fundamentally new implementation (see Appendix 7.9 for a detailed FLOP comparison).
@@ 5 Conclusion
ButterflyQuant demonstrates that continuous parameterization of orthogonal transforms fundamentally changes what is achievable in extreme quantization. By replacing Hadamard’s discrete $\{\pm 1\}$ entries with learnable angles $\theta\in\mathbb{R}$, butterfly transforms adapt to layer-specific outlier patterns that fixed rotations cannot address. This simple but powerful insight enables practical 2-bit quantization that consistently outperforms existing rotation-based methods. Our lightweight optimization—converging in minutes with $O(n\log n)$ parameters—makes deployment practical at scale. Moreover, since the learned butterfly transform preserves the same computational graph as the fast Hadamard transform, deployment can leverage existing optimized CUDA kernels with only minor modifications to accommodate pre-computed twiddle factors. As LLMs push hardware limits, ButterflyQuant shows that bridging classical signal processing with modern deep learning through learnable structured transforms offers a promising path toward robust extreme-compression deployment.
Impact Statement
This work contributes to democratizing access to large language models by enabling their deployment on consumer-grade hardware, potentially reducing the resource barriers that currently concentrate AI capabilities among well-funded entities. By compressing models like LLaMA-70B from 140GB to approximately 17.5GB with minimal accuracy loss, our method also reduces the energy footprint of inference, contributing to more sustainable AI deployment at scale. However, we acknowledge that lowering deployment barriers may inadvertently facilitate misuse of powerful language models by actors with limited resources for safety measures. Importantly, our technique does not introduce new model capabilities or training methods—it strictly addresses model compression—and thus the ethical considerations remain those inherent to the underlying pre-trained models being quantized. We encourage practitioners deploying quantized models to maintain the same safety protocols and usage guidelines as the original full-precision versions.